OpenStudy (anonymous):

What is x in the equation: 6x + 20 = -2x^2 ?

OpenStudy (anonymous):

5,-2

OpenStudy (anonymous):

my bad. I ment 17, not 20

OpenStudy (anonymous):

-17(+/-) sqrt(17^2-4(6)(2)) can you do that

OpenStudy (anonymous):

and all over 2(2)

OpenStudy (anonymous):

-17 +/- sqrt(241).. the rest I'm not sure about

OpenStudy (anonymous):

sorry I have led you incorrectly

OpenStudy (anonymous):

I will do the problem and show the steps,

OpenStudy (anonymous):

\[6x + 20 = -2x^2 \implies 2x^2+6x+20=0 \implies no~real~solution\]

OpenStudy (anonymous):

I say no real solution since the determinant is negative.

OpenStudy (anonymous):

yep so -6+/- sqrt(36-5(17)(2))/4 x= (-3-5i)/2 x=(-3+5i)/2

OpenStudy (anonymous):

i ment 17, not 20 @AnimalAin

OpenStudy (anonymous):

OK, but my answer is still the same...

OpenStudy (anonymous):

That would be the discriminant, not the determinant.... foolishness....

OpenStudy (anonymous):

medal please